OpenStudy (danielbarriosr1):

Please help me creating 2 equation

OpenStudy (danielbarriosr1):

Carol has three times as many 10-dollar bills as 5-dollar bills. She has a total of $350. How many 5-dollar bills does Carol have?

OpenStudy (danielbarriosr1):

@mathmale

OpenStudy (danielbarriosr1):

@primeralph

OpenStudy (anonymous):

Let's say x=10-dollar bills y=5-dollar bills So 10 dollar bills=3 * 5-dollar bills so x=3y So find the total of each you have 10-dollar bills + 5-dollar bills = $350 So x+y=350

OpenStudy (danielbarriosr1):

and the other equation?

OpenStudy (anonymous):

I gave 2. x=3y and x+y=350 Now you can solve.

OpenStudy (danielbarriosr1):

Oh so x=3y Got it

OpenStudy (danielbarriosr1):

x=525/2 y=175/2

OpenStudy (danielbarriosr1):

But possible answers are: A. 10 B. 14 C. 30 D. 42

OpenStudy (danielbarriosr1):

Are you sure those are the correct equations?

OpenStudy (danielbarriosr1):

@MrsDavidRandle

OpenStudy (anonymous):

Yes I am.

OpenStudy (danielbarriosr1):

But what is the answer then?

OpenStudy (anonymous):

Actually no those can't be correct. x=3y 10x+5y=350 Ok sorry!

OpenStudy (anonymous):

Those should get you the answer! Sorry my num lock wasn't on so it type them the first time hahaa.

OpenStudy (danielbarriosr1):

Y=10

OpenStudy (anonymous):

That's what I got too.